# NyayaRAG: Realistic Legal Judgment Prediction with RAG under the Indian Common Law System

### v1

@@ Abstract
Shubham Kumar Nigam1∗†   Balaramamahanthi Deepak Patnaik1∗ Shivam Mishra1∗ Ajay Varghese Thomas2∗Noel Shallum3Kripabandhu Ghosh4Arnab Bhattacharya1 1 IIT Kanpur, India  2 SRM Institute of Science and Technology, India 3 IISER Kolkata, India  4 Symbiosis Law School Pune, India {sknigam, deepak, shivammishra, arnabb}@cse.iitk.ac.in at4231@srmist.edu.inkripaghosh@iiserkol.ac.in,noelshallum@gmail.com
\$*\$\$*\$footnotetext: These authors contributed equally to this work\$$\dagger$\$$\dagger$\$footnotetext: Corresponding author
@@ 1 Introduction
Previous studies such as Malik et al. (2021); Nigam et al. (2024b, 2025a) have focused on predicting legal outcomes using the current case document, including sections like facts, arguments, issues, reasoning, and decision. More recent efforts have narrowed the scope to factual inputs alone (Nigam et al., 2024a, 2025b), yet these systems still operate in a vacuum, without considering how courts naturally rely on applicable laws and prior rulings. In reality, judges rarely decide in isolation; instead, they actively refer to relevant precedent and statutory law. To bridge this gap, we propose a framework that more closely mirrors actual courtroom conditions by explicitly incorporating external legal knowledge during inference.
@@
Our work moves beyond fact-only or self-contained models by replicating a more faithful legal reasoning pipeline aligned with Indian jurisprudence. We hope that NyayaRAG opens new directions for building interpretable, retrieval-aware AI systems in legal settings, particularly in resource-constrained yet precedent-driven judicial systems like India’s. For the sake of reproducibility, we have made our dataset, code, and RAG-based pipeline implementation via a GitHub repository [^1] https://github.com/ShubhamKumarNigam/RAGLegal.
@@ 2 Related Work
Recent advancements in natural language processing (NLP) and large language models (LLMs) have significantly improved the performance of question answering (QA) and legal decision support systems. Transformer-based architectures such as BERT (Devlin et al., 2018), GPT (Radford et al., 2019), and their instruction-tuned successors have led to robust capabilities in knowledge-intensive and multi-hop reasoning tasks. The integration of external information via Retrieval-Augmented Generation (RAG) has emerged as a particularly effective approach for enhancing generation fidelity and reducing hallucinations (Han et al., 2024; Hei et al., 2024).
Within the legal domain, Legal Judgment Prediction (LJP) has seen significant progress, with models trained to infer outcomes based on factual and procedural components of court cases (Strickson and De La Iglesia, 2020; Xu et al., 2020; Feng et al., 2023). In the Indian legal context, the ILDC corpus (Malik et al., 2021) and its extended variants (Nigam et al., 2024b; Nigam and Deroy, 2023) have enabled the development of supervised and instruction-tuned models for both judgment prediction and explanation. The emergence of domain-specific datasets and architectures has allowed LJP systems to move from simple binary classification to more complex reasoning tasks aligned with real judicial behavior (Vats et al., 2023).
Parallel to these developments, there has been a sharp rise in interest in RAG techniques for legal NLP. Several benchmark and system-level contributions have explored how retrieval-enhanced generation can be leveraged to assist legal professionals, improve legal QA systems, and support document analysis. Notably, LegalBench-RAG (Pipitone and Alami, 2024) introduced a benchmark suite for evaluating RAG in the legal domain. Survey papers like (Hindi et al., 2025) provide comprehensive overviews of techniques aimed at improving RAG performance, factual grounding, and interpretability in legal settings.
Several system-level contributions have demonstrated the power of RAG in specialized applications. Graph-RAG for Legal Norms (de Martim, 2025) and Bridging Legal Knowledge and AI (Barron et al., 2025) proposed methods to integrate structured legal knowledge such as statutes and normative hierarchies into the retrieval pipeline. Similarly, CBR-RAG (Wiratunga et al., 2024) applied case-based reasoning to leverage historical decisions, showing strong gains in legal question answering. HyPA-RAG (Kalra et al., 2024) explored hybrid parameter-adaptive retrieval to dynamically adjust context based on query specificity.
Further domain-specific applications include AI-powered legal assistants like Legal Query RAG (Wahidur et al., 2025) and RAG-based solutions for dispute resolution in housing law (Rafat, 2024). Optimizing Legal Information Access (Amato et al., 2024) showcased federated RAG architectures for secure document retrieval, and Augmenting Legal Judgment Prediction with Contrastive Case Relations (Liu et al., 2022) illustrated the benefits of encoding contrastive precedents for predictive reasoning.
@@ 3 Task Description
[…]
@@ 4.2.1 Case Text
Each judgment includes complete narrative content such as factual background, party arguments, legal issues, reasoning, and verdict. Due to length constraints exceeding model context windows, we summarized these documents using Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024), which supports up to 32k tokens. The summarization preserved critical legal elements through carefully designed prompts (see Table 2).
@@ 4.2.4 Previous Similar Cases
[…]
@@ 5.3 Inference Setup
We use the LLaMA-3.1 8B Instruct Dubey et al. (2024) model for all experiments in a few-shot prompting setup. Each input sequence, composed according to one of the pipeline templates, is paired with a relevant prompt. The model is required to output:
@@ 6 Evaluation Metrics
    Lexical-based Evaluation: We utilized standard lexical similarity metrics, including Rouge-L Lin (2004), BLEU Papineni et al. (2002), and METEOR Banerjee and Lavie (2005). These metrics measure the overlap and order of words between the generated explanations and the reference texts, providing a quantitative assessment of the lexical accuracy of the model outputs.
@@
    Semantic Similarity-based Evaluation: To capture the semantic quality of the generated explanations, we employed BERTScore Zhang et al. (2020), which measures the semantic similarity between the generated text and the reference explanations. Additionally, we used BLANC Vasilyev et al. (2020), a metric that estimates the quality of generated text without a gold standard, to evaluate the model’s ability to produce semantically meaningful and contextually relevant explanations.
@@
    LLM-based Evaluation (LLM-as-a-Judge): To complement traditional metrics, we incorporate an automatic evaluation strategy that uses large language models themselves as evaluators, commonly referred to as LLM-as-a-Judge. This evaluation is crucial for assessing structured argumentation and legal correctness in a format aligned with expert judicial reasoning. We adopt G-Eval Liu et al. (2023), a GPT-4-based evaluation framework tailored for natural language generation tasks. G-Eval leverages chain-of-thought prompting and structured scoring to assess explanations along three key criteria: factual accuracy, completeness & coverage, and clarity & coherence. Each generated legal explanation is scored on a scale from 1 to 10 based on how well it aligns with the expected content and a reference document. The exact prompt format used for evaluation is shown in Appendix Table 9. For our experiments, we use the GPT-4o-mini model to generate reliable scores without manual intervention. This setup provides an interpretable, unified judgment metric that captures legal soundness, completeness of reasoning, and logical coherence, beyond what traditional similarity-based metrics can offer.
@@
    Inter-Annotator Agreement (IAA): To ensure the reliability and consistency of expert judgments, we compute standard IAA statistics, including Fleiss’ Kappa, Cohen’s Kappa, Krippendorff’s Alpha, Intraclass Correlation Coefficient (ICC), and Pearson Correlation. These metrics quantify the degree of agreement across expert raters, reinforcing the credibility of the expert evaluation framework. Full details and scores are available in Appendix B.
7 Results and Analysis
@@ 7.2 Explanation Generation Quality
To complement automatic evaluations, we also conducted a small-scale expert evaluation involving experienced legal professionals. Each expert independently rated a subset of model-generated explanations based on factual accuracy, legal relevance, and completeness using a 10-point Likert scale. The results from this human evaluation corroborated the trends observed in automatic metrics. Notably, the CaseText + Statutes pipeline received the highest expert score among all configurations, reinforcing the positive impact of statutory knowledge on explanation quality. In contrast, fact-only pipelines again received the lowest expert ratings, echoing concerns about their insufficient legal reasoning depth.
To ensure the reliability of expert scores, we conducted a detailed Inter-Annotator Agreement (IAA) analysis across multiple evaluation dimensions. The IAA results (Appendix B, Table 5) reveal substantial agreement between legal experts, with consistently high values across Fleiss’ Kappa, ICC, and Krippendorff’s Alpha. These findings reinforce the consistency and trustworthiness of our expert-based human evaluation framework.
Overall, the results emphasize the effectiveness of Retrieval-Augmented Generation (RAG) when paired with structured legal content, especially statutes, in producing accurate, interpretable, and legally coherent explanations. The inclusion of G-Eval and expert ratings provides a multifaceted lens for assessing explanation quality, bridging the gap between automatic evaluation and real-world legal judgment standards.
@@ A.2 Judgment Prediction Hyper-parameters
Appendix B Inter-Annotator Agreement (IAA) for Expert Evaluation
B.1 IAA Metrics and Methodology
To ensure the reliability and consistency of expert ratings on the quality of generated legal explanations, we computed five widely accepted inter-rater agreement metrics:
  • Fleiss’ Kappa Fleiss (1971): Evaluates agreement among multiple raters for categorical judgments, adjusting for chance.
  • Cohen’s Kappa Cohen (1960): Measures the pairwise agreement between two annotators, controlling for expected chance agreement.
  • Intraclass Correlation Coefficient (ICC) Shrout and Fleiss (1979): Assesses the degree of consistency in continuous ratings across multiple raters.
  • Krippendorff’s Alpha Krippendorff (2018): A versatile metric capable of handling varying scales and missing data, suitable for ordinal and interval data.
  • Pearson Correlation Coefficient Benesty et al. (2009): Quantifies the strength of the linear relationship between expert rating scores.
Three experienced legal experts independently rated a shared subset of model-generated legal explanations on a 10-point Likert scale, considering factual accuracy, legal relevance, and completeness. The raters were blind to the model configurations to avoid bias and promote objective assessment.
B.2 IAA Findings and Observations
Interpretation:
The overall inter-annotator agreement across all evaluation settings demonstrates moderate to substantial reliability. In the Single partition, pipelines such as CaseText + Statutes and CaseText + Statutes + Precedents achieved the highest agreement scores across most metrics (e.g., Fleiss’ $\kappa>$ 0.49, ICC almost 0.60), indicating stronger consensus among experts regarding their quality. This is aligned with the higher expert scores and other automatic evaluation metrics for these pipelines.
In contrast, pipelines using only factual input or combining facts with retrieved statutes and precedents (CaseFacts Only and Facts + Statutes + Precedents) yielded relatively lower agreement scores (e.g., Fleiss’ $\kappa<0.35$, ICC <0.50<0.50< 0.50), reflecting the increased ambiguity or inconsistency in explanation quality when limited or noisy contextual information is used.
The Multi partition exhibits slightly lower agreement metrics overall, potentially due to the complexity introduced by multiple judgment labels per case. Still, pipelines with richer legal context (CaseText + Statutes, CaseText + Statutes + Precedents) maintained comparatively higher consistency among annotators.
Conclusion:
These results reinforce the interpretability and credibility of our expert evaluation process. The observed agreement levels validate that the rating protocol is sufficiently robust to distinguish between explanation quality across different legal input pipelines. Moreover, the findings corroborate trends observed through both automatic and LLM-based evaluation metrics.
Single Partition
CaseText Only 0.42 0.47 0.55 0.49 0.58
CaseText + Statutes 0.51 0.55 0.61 0.57 0.65
CaseText + Precedents 0.37 0.42 0.50 0.45 0.52
CaseText + Previous Similar Cases 0.41 0.45 0.54 0.47 0.56
CaseText + Statutes + Precedents 0.49 0.52 0.59 0.54 0.62
CaseFacts Only 0.34 0.39 0.47 0.43 0.49
Facts + Statutes + Precedents 0.29 0.34 0.42 0.38 0.44
Multi Partition
CaseText Only 0.40 0.45 0.53 0.48 0.57
CaseText + Statutes 0.50 0.54 0.60 0.56 0.64
CaseText + Precedents 0.36 0.41 0.49 0.44 0.51
CaseText + Previous Similar Cases 0.39 0.44 0.52 0.46 0.55
CaseText + Statutes + Precedents 0.47 0.51 0.58 0.52 0.60
CaseFacts Only 0.32 0.37 0.46 0.41 0.48
Facts + Statutes + Precedents 0.28 0.33 0.41 0.36 0.43

### v2

@@ Abstract
Shubham Kumar Nigam1,5∗†  Balaramamahanthi Deepak Patnaik1∗  Shivam Mishra1∗ Ajay Varghese Thomas2∗Noel Shallum3Kripabandhu Ghosh4Arnab Bhattacharya1 1IIT Kanpur, India,  2SRM Institute of Science and Technology, India 3Symbiosis Law School Pune, India  4IISER Kolkata, India 5University of Birmingham, Dubai, United Arab Emirates {shubhamkumarnigam, bdeepakpatnaik2002, shivam1602m, ajaythomas.work, noelshallum} @gmail.com kripaghosh@iiserkol.ac.in,arnabb@cse.iitk.ac.in
[…]
@@ 1 Introduction
Previous studies such as malik-etal-2021-ildc; nigam-etal-2024-legal; nigam-etal-2025-nyayaanumana have focused on predicting legal outcomes using the current case document, including sections like facts, arguments, issues, reasoning, and decision. More recent efforts have narrowed the scope to factual inputs alone (nigam-etal-2024-rethinking; FactLegalLlama), yet these systems still operate in a vacuum, without considering how courts naturally rely on applicable laws and prior rulings. In reality, judges rarely decide in isolation; instead, they actively refer to relevant precedent and statutory law. To bridge this gap, we propose a framework that more closely mirrors actual courtroom conditions by explicitly incorporating external legal knowledge during inference.
@@
Our work moves beyond fact-only or self-contained models by replicating a more faithful legal reasoning pipeline aligned with Indian jurisprudence. We hope that NyayaRAG opens new directions for building interpretable, retrieval-aware AI systems in legal settings, particularly in resource-constrained yet precedent-driven judicial systems like India’s. For the sake of reproducibility, we have implemented our data set, code, and RAG-based pipeline through a GitHub repository [^1] https://github.com/ShubhamKumarNigam/NyayaRAG.
@@ 2 Related Work
Recent advances in natural language processing (NLP) and large language models (LLMs) have transformed knowledge-intensive tasks, including question answering and decision support. Transformer-based architectures such as BERT (devlin2018bert), GPT (radford2019language), and their instruction-tuned successors demonstrate strong multi-hop reasoning and contextual generation abilities. Retrieval-Augmented Generation (RAG) further improves factual grounding and reduces hallucinations by combining external document retrieval with generative models (han2024rag; hei2024dr).
The domain of Legal Judgment Prediction (LJP) has evolved from outcome classification toward explanation generation and realistic reasoning. Foundational work explored case outcome prediction in ECHR and Chinese datasets (aletras2016predicting; chalkidis2019neural; xiao2018cail2018), inspiring subsequent benchmarks such as CAIL2018 and ECHR-CASES. Recent approaches have incorporated statutes, reasoning structures, and event-based features (feng-etal-2022-legal; feng2023criminal).
Within the Indian legal ecosystem, the ILDC corpus (malik-etal-2021-ildc) laid the foundation for Court Judgment Prediction and Explanation (CJPE). This has been expanded in PredEx (nigam2024legaljudgmentreimaginedpredex), NyayaAnumana (nigam2024nyayaanumana), LegalSeg (nigam2025legalseg), TathyaNyaya (nigam2025tathyanyaya), and IBPS for bail prediction (srivastava2025ibps), which collectively focus on segmentation, factual reasoning, and judgment outcome explanation. Efforts such as nigam2024rethinking and 10.1145/3632754.3632765 emphasize fact-based LJP as a realistic alternative to full-text reasoning, aligning predictions more closely with early-phase decision-making. Other works explore semantic segmentation (malik2021semantic), legal QA (nigam2023legal), and structured document generation (nigam2025structured). These resources and models highlight the significance of domain-specific adaptation for Indian law, an underrepresented common law system.
Precedent integration is central to common law reasoning. Early precedent-aware models such as zhao-etal-2018-learning focused on Chinese law, while recent works like PLJP (wu2023precedent), LexKeyPlan (santosh2025lexkeyplan), and precedent-enhanced ECHR prediction (santosh2024incorporating) examine retrieval of historical judgments to guide predictions. These align with broader multilingual and cross-jurisdictional LJP studies (niklaus2021swiss; kapoor-etal-2022-hldc). While these efforts highlight the role of past cases, their settings differ significantly from India’s unique common law context where both statutory law and judicial precedents guide reasoning.
Retrieval-augmented pipelines are increasingly applied to legal NLP tasks. LegalBench-RAG (pipitone2024legalbench) and CLERC (hou2024clerc) introduced benchmarks for retrieval-augmented legal reasoning. Specialized systems like CBR-RAG (wiratunga2024cbr), Graph-RAG for norms (de2025graph), and LexKeyPlan (santosh2025lexkeyplan) explore structured retrieval, case-based reasoning, and keyphrase-driven planning. Applications range from legal assistants and dispute resolution (rafat2024ai; 10887211) to federated secure access architectures (amato2024optimizing). These works demonstrate the value of augmenting generative reasoning with statutory and precedent-based retrieval but primarily address civil law or international corpora.
Our work contributes to this line by contextualizing RAG for the Indian common law system, simulating how judges reason with case facts, statutory provisions, and precedents. While prior RAG frameworks in law (wiratunga2024cbr; hou2024clerc; wu2023precedent; santosh2024incorporating) establish the feasibility of retrieval-augmented pipelines, they have not been applied to Indian courts. By integrating factual, statutory, and precedential inputs in a unified framework, NyayaRAG extends the landscape of LJP toward realistic, explainable, and domain-grounded prediction and explanation.
@@ 3 Task Description
Prediction Task: Based on the summarized factual description $D$ and the retrieved top-$k$ (e.g., $k=3$) similar legal documents (statutes or precedents), the model predicts the likely court judgment. The prediction label $y\in\{0,1\}$ indicates whether the appeal is fully rejected (0) or fully/partially accepted (1). This binary framing captures the most common forms of judicial decisions in Indian appellate courts.
@@ 4.1 Dataset Compilation
4.2 Single vs. Multi Partitioning
Another challenge in the SCI proceedings is the presence of cases containing multiple petitions under a single case identifier, where each petition may lead to distinct outcomes. To handle this, we partition the dataset into two subsets. The single partition represents binary classification tasks (accept/reject) and includes cases with a single petition or multiple petitions that share the same outcome. The multi partition, a superset of single, contains cases with multiple petitions leading to different decisions, introducing a more complex multi-label prediction setting (e.g., partially accepted outcomes). For these cases, if even one appeal within the group is accepted, the case is assigned the label accepted. Predicting such mixed outcomes remains challenging and will be further explored in future work.
@@ 4.3.1 Case Text
Each judgment includes complete narrative content such as factual background, party arguments, legal issues, reasoning, and verdict. Due to length constraints exceeding model context windows, we summarized these documents using Mixtral-8x7B-Instruct-v0.1 (Jiang2024MixtralOE), which supports up to 32k tokens. The summarization preserved critical legal elements through carefully designed prompts (see Table 2).
@@ 4.3.4 Previous Similar Cases
  • Top-$k$ Retrieval: Using ChromaDB, we retrieved the top-3 most semantically similar cases for each document based on cosine similarity.
@@ 5.3 Inference Setup
We use the LLaMA-3.1 8B Instruct dubey2024llama model for all experiments in a few-shot prompting setup. Each input sequence, composed according to one of the pipeline templates, is paired with a relevant prompt. The model is required to output:
@@ 6 Evaluation Metrics
    Lexical-based Evaluation: We utilized standard lexical similarity metrics, including Rouge-L lin-2004-rouge, BLEU papineni-etal-2002-bleu, and METEOR banerjee-lavie-2005-meteor. These metrics measure the overlap and order of words between the generated explanations and the reference texts, providing a quantitative assessment of the lexical accuracy of the model outputs.
@@
    Semantic Similarity-based Evaluation: To capture the semantic quality of the generated explanations, we employed BERTScore BERTScore, which measures the semantic similarity between the generated text and the reference explanations. Additionally, we used BLANC blanc, a metric that estimates the quality of generated text without a gold standard, to evaluate the model’s ability to produce semantically meaningful and contextually relevant explanations.
@@
    LLM-based Evaluation (LLM-as-a-Judge): To complement traditional metrics, we incorporate an automatic evaluation strategy that uses large language models themselves as evaluators, commonly referred to as LLM-as-a-Judge. This evaluation is crucial for assessing structured argumentation and legal correctness in a format aligned with expert judicial reasoning. We adopt G-Eval liu-etal-2023-g, a GPT-4-based evaluation framework tailored for natural language generation tasks. G-Eval leverages chain-of-thought prompting and structured scoring to assess explanations along three key criteria: factual accuracy, completeness & coverage, and clarity & coherence. Each generated legal explanation is scored on a scale from 1 to 10 based on how well it aligns with the expected content and a reference document. The exact prompt format used for evaluation is shown in Appendix Table 9. For our experiments, we use the GPT-4o-mini model to generate reliable scores without manual intervention. This setup provides an interpretable, unified judgment metric that captures legal soundness, completeness of reasoning, and logical coherence, beyond what traditional similarity-based metrics can offer.
@@
    Inter-Annotator Agreement (IAA): To further ensure the reliability and consistency of expert judgments, we computed multiple standard inter-rater agreement metrics. Specifically, Fleiss’ Kappa fleiss1971measuring was used to evaluate the overall agreement among multiple raters, Cohen’s Kappa cohen1960coefficient captured pairwise agreement while adjusting for chance, and the Intraclass Correlation Coefficient (ICC) shrout1979intraclass measured the reliability of continuous ratings across raters. In addition, Krippendorff’s Alpha krippendorff2018content provided a robust measure suitable for ordinal scales and missing data, while the Pearson Correlation Coefficient benesty2009pearson quantified the linear consistency of expert scores. The results demonstrated substantial agreement across these different measures, reinforcing the credibility of the evaluation and lending strong support to the robustness of our findings.
Together, these metrics provide a comprehensive, multi-perspective evaluation of our system’s ability to both predict judicial outcomes and generate legally coherent, interpretable rationales grounded in retrieved context.
@@ 7.2 Explanation Generation Quality
The results of this evaluation aligned closely with the trends observed in automatic metrics (Table 4). Pipelines enriched with statutory provisions, particularly CaseText + Statutes, consistently received the highest expert ratings, highlighting the value of legal grounding in improving explanation quality. In contrast, pipelines relying solely on factual input, such as CaseFacts Only, obtained the lowest expert scores, underscoring their lack of interpretive depth.
To assess the reliability of these expert judgments, we performed a detailed Inter-Annotator Agreement (IAA) analysis across multiple evaluation dimensions, Table 5. The IAA results revealed moderate to substantial agreement among annotators. For the Single partition, pipelines such as CaseText + Statutes and CaseText + Statutes + Precedents achieved higher consistency (Fleiss’ $\kappa>0.49$, ICC $\approx 0.60$), reflecting strong consensus on their explanation quality. In contrast, fact-only or noisy-input pipelines showed weaker agreement (Fleiss’ $\kappa<0.35$, ICC <0.50<0.50), suggesting greater variability in expert perception of their outputs.
The Multi partition exhibited slightly lower agreement overall, likely due to the additional complexity of multi-label judgments. Nevertheless, richer legal context pipelines again demonstrated higher consistency among experts compared to fact-only inputs. These findings not only validate the robustness of our expert evaluation protocol but also corroborate trends from lexical, semantic, and LLM-based evaluations.
Taken together, the results emphasize that Retrieval-Augmented Generation, when paired with structured legal inputs such as statutes and precedents, produces explanations that are more accurate, interpretable, and legally coherent. The combination of automatic metrics, LLM-based judgment, and human expert ratings provides a multifaceted and credible framework for assessing explanation quality in legal judgment prediction.
Single Partition
CaseText Only 0.42 0.47 0.55 0.49 0.58
CaseText + Statutes 0.51 0.55 0.61 0.57 0.65
CaseText + Precedents 0.37 0.42 0.50 0.45 0.52
CaseText + Previous Similar Cases 0.41 0.45 0.54 0.47 0.56
CaseText + Statutes + Precedents 0.49 0.52 0.59 0.54 0.62
CaseFacts Only 0.34 0.39 0.47 0.43 0.49
Facts + Statutes + Precedents 0.29 0.34 0.42 0.38 0.44
Multi Partition
CaseText Only 0.40 0.45 0.53 0.48 0.57
CaseText + Statutes 0.50 0.54 0.60 0.56 0.64
CaseText + Precedents 0.36 0.41 0.49 0.44 0.51
CaseText + Previous Similar Cases 0.39 0.44 0.52 0.46 0.55
CaseText + Statutes + Precedents 0.47 0.51 0.58 0.52 0.60
CaseFacts Only 0.32 0.37 0.46 0.41 0.48
Facts + Statutes + Precedents 0.28 0.33 0.41 0.36 0.43
@@ 9 Conclusion and Future Scope
Acknowledgements
We would like to express our sincere gratitude to the anonymous reviewers for their constructive feedback and valuable suggestions, which greatly improved the quality of this paper. We also thank the student research assistants and legal experts from various law colleges for their dedicated efforts in annotation and expert evaluation. Their contributions were instrumental in ensuring the quality and reliability of the dataset and evaluation process. We gratefully acknowledge the support of BharatGen, India, for providing access to computational resources and hardware infrastructure used in this research. Their assistance played a vital role in model training and large-scale experimentation. The majority of this work was conducted while the first author was affiliated with the Indian Institute of Technology Kanpur. The author is currently affiliated with the University of Birmingham Dubai.